A 12.45 kg stone falls from the top of a cliff and strikes the ground at a speed of 46.2 m/s. What is the height of the cliff ?


1
Expert's answer
2021-03-08T08:28:09-0500
"v^2=2gh\\46.2^2=2(9.8)h\\h=108.9\ m"